Jordan Dance Company is considering whether to sell his dance studio to Sophia Dance Company.  Jordan purchased her dance studio three years ago at a cost of $40,000.  At that time, she anticipated selling the studio after seven years for $5,000.  Her dance studio holds 400 students and she earns $250 per student (paid at the end of the year).  Jordan will retire from dancing today if she sells, otherwise she will continue teaching for the original seven years she had planned when she opened her studio (i.e., another four years).

Required:  Using time-value-of-money calculations, what would be the fair price for Sophia to pay Jordan?  Assume a tax rate of 30%, straight-line depreciation [the same for financial reporting or taxes; i.e., she allocates the depreciable basis (historical cost – salvage value) over the time period, rather than the historical cost of the asset], and an annual discount rate is six-percent.   
Answer #1
years cash flows discount rate @6% discounted cash flows
0 (40000) 1 (40000)
1 400*250=100000 1/1.06=0.94 94000
2 100000 0.89 89000
3 100000 0.84 84000
4 100000 0.79 79000
5 100000 0.75 75000
6 100000 0.70 70000
7 100000 0.66 66000
total income 517000

total discounted income = $517000

less :tax @30% = $ 517000*30%=$155100

net income = $361900

Fair price for sophia to jordan is $ 361900